Observers: Linhua Jiang, Emily McLinden

6:40 pm: start luci.

7:25 pm: slew to Emily's target obj1, and do acquisition.

7:47 pm: start observing obj1 (clear sky and good seeing ~0.8").

8:55 pm: finish obj1, and take calibration for obj1.

9:00 pm: slew to Emily's target obj2, and do acquisition.

9:23 pm: start observing obj2 (clear sky and good seeing).

10:40 pm: finish obj2, and take calibration for obj2.

10:50 pm: slew to SDF instead of z~9 galaxy. We have to figure out the reason of positioning failure last night. We try to use N1.8 for acquisition instead of N3.75. We still have a big offset between software predicted positions and really observe object positions. We manually do offsets (the good thing is that there are a few bright objects in the slits) a few times until all bright objects fit into slits. A painful procedure.

12:10 am: we start first science exposure. We now confirmed that the failure of positioning of last night was due to the fact that control software was not updated for new camera.

12:30 am: we get first science image. We get all objects expected in the image, meaning positioning was good. However, there are serious persistence from images taken 1.5 hours ago. This is something you cannot get rid of.

3:00 am: keep observing, flexure compensation not working well.

4:00 am: finished SDF, take calibration (it's very windy now).

4:15 am: very windy, check wind speed by pointing telescope to right direction.

4:20 am: it's ok, slew to Benjamin's target.

5:10 am: finished two 15min exposure and take calibration.

5:20 am: done.

-- XiaohuiFan - 18 Apr 2012
Topic revision: r8 - 19 Apr 2012, XiaohuiFan
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback